The area of a playground is 126 yards the width of the playground is 5 yards longer then it's like find the length and width of the playground
Answer by Cromlix(4381)

Hi there,
Make length = x
Width = x + 5
Area = Width * length (* times)
126 = (x + 5) * x
126 = x^2 + 5x
Sort into ax^2 + bx + c form
x^2 + 5x - 126 = 0
(x - 9)(x + 16) = 0
x + 16 = 0
x + -16 (disregard as -ve)
x - 9 = 0
x = 9
Length = 9yds
Width = 14yds
Hope this helps :-)